jagr2713 (jagr2713):

Trick Question: A + D + 8 = ?

jagr2713 (jagr2713):

@Molly_Qian give other people a chance

OpenStudy (anonymous):

ughh ok...

jagr2713 (jagr2713):

@pooja195 @Zale101

jagr2713 (jagr2713):

@bahrom7893

OpenStudy (compassionate):

A + D + H + D + 8 = too many kids to handle

jagr2713 (jagr2713):

Lol Cmon some can get this :D

jagr2713 (jagr2713):

You guys want a hint?

OpenStudy (anonymous):

Yes lol

jagr2713 (jagr2713):

@TheSmartOne

OpenStudy (anonymous):

am i allowed to take an educated guess?

jagr2713 (jagr2713):

sure

OpenStudy (anonymous):

a date?

jagr2713 (jagr2713):

nope

OpenStudy (anonymous):

good.

jagr2713 (jagr2713):

@pooja195 got this :D or @Luigi0210

jagr2713 (jagr2713):

ok no one is getting this so let me give a hint lol i am going to wait another 20secs

TheSmartOne (thesmartone):

Waiiiit

jagr2713 (jagr2713):

oK :d

jagr2713 (jagr2713):

you got this

TheSmartOne (thesmartone):

\(\sf A + D + 8 = D + A + 8 = 8 + A + D = A + 8 + D = D + 8 + A\)

jagr2713 (jagr2713):

lol ok Hint time Answer as fast as u can

OpenStudy (anonymous):

plz close

TheSmartOne (thesmartone):

8 A.D ?

jagr2713 (jagr2713):

Ok everyone say A D FASTTTT and tell me what u get

OpenStudy (anonymous):

this is annoying

Nnesha (nnesha):